Cabrina and Dabney are attending a conference. After the​ conference, Cabrina drives home to Boise at an average speed of 75 miles per hour and Dabney drives home to Portland at an average speed of 60 miles per hour. If the sum of their driving times is 10.1 hours and if the sum of the distances driven is 711 ​miles, determine the time each woman spent driving home.
Answer by ankor@dixie-net.com(22740)

Cabrina and Dabney are attending a conference.
After the​ conference, Cabrina drives home to Boise at an average speed of 75 miles per hour and Dabney drives home to Portland at an average speed of 60 miles per hour.
If the sum of their driving times is 10.1 hours and if the sum of the distances driven is 711 ​miles, determine the time each woman spent driving home.
:
let t = C's driving time
the total time is 10.1 hrs, therefore
(10.1-t) = D's driving time
:
Write a dist equation, dist. speed * time
C's dist + D's dist = 711 mi
60t + 75(10.1-t) = 711
60t + 757.5 - 75t = 711
60t - 75t = 711 - 757.5
-15t = -46.5
t = -46.5/-15
t = 3.1 hrs is C's travel time
then
10.1-3.1 = 7 hr is D's travel time
:
:
Check this by finding the distances each traveled
60*3.1 = 186 mi
75 * 7 = 525 mi
-----------------
total dis 711 mi